James needs to mix a 10% alcohol solution with a 40% alcohol solution to create 100 milliliters of a 37% solution. How many milliliters of each solution must James use?
Answer by richwmiller(17219)

We need a total of 100 mL at 37%
We want to know how many mL at 10% to mix with 40%
a+b=100,
0.1*a+0.4*b=0.37*100
a=100-b
0.1*(100-b)+0.4*b=37
10-0.1b+0.4*b=37
0.3*b=27
b=90
a=100-b
a=10 mL at 10%
b=90 mL at 40%
check
0.1*10+0.4*90=0.37*100
1+36=37
37=37
ok
codemx1